# Descent Methods for Vector Optimization Problems: A Majorization-Minimization Perspective

### v2

@@ Abstract
In this paper, we develop a unified framework and convergence analysis of first-order methods for vector optimization problems (VOPs) from a majorization-minimization perspective. By selecting different surrogate functions, the general method can be reduced to various existing first-order methods, both with and without line search. Our unified convergence analysis reveals that the slow convergence of the steepest descent method is primarily attributed to the significant gap between the surrogate and objective functions. Consequently, narrowing this surrogate gap can enhance the performance of first-order methods. Interestingly, we elucidate that selecting a tighter surrogate function is equivalent to using an appropriate base of the dual cone in the direction-finding subproblem. Building on this insight, we employ the Barzilai-Borwein method to narrow the surrogate gap and propose a Barzilai-Borwein descent method for VOPs (BBDVO) with polyhedral cones. By reformulating the subproblem, we provide a novel perspective on the Barzilai-Borwein descent method, bridging the gap between this method and the steepest descent method. Furthermore, by selecting an appropriate base and reformulating the subproblem in the Newton method, we also provide a novel perspective on the preconditioned Barzilai-Borwein method. Finally, several numerical experiments are presented to validate the efficiency of the BBDVO.
@@ 1 Introduction
This paper focuses on the following vector optimization problem:
@@
Let $K^{*}=\{c^{*}\in\mathbb{R}^{m}:\langle c^{*},y\rangle\geq 0,\forall y\in K\}$ be the positive polar cone of $K$, and $C$ be a compact convex set such that $0\notin C$ and $\mathrm{cone}(C)=K^{*}$, namely, $C$ is a base of $K^{*}$. In vector optimization, it is often impossible to improve all objectives simultaneously with respect to the partial order. Therefore, the concept of optimality is defined as efficiency [25], meaning that there is no better solution for an efficient solution. Specifically, the problem (VOP) corresponds to a multiobjective optimization problem when $K=\mathbb{R}^{m}_{+}$, where +$\mathbb{R}^{m}_{+}$ denotes the non-negative orthant of $\mathbb{R}^{m}$. Various applications of multiobjective optimization problems (MOPs) can be found in engineering [35], economics [45, 20], management science [15], environmental analysis [32], machine learning [41, 47], etc. Although many real-world problems reformulated as vector-valued problems adhere to the partial order induced by +$\mathbb{R}^{m}_{+}$, some applications, such as portfolio selection in securities markets [1, 2], require partial orders induced by closed convex cones other than the non-negative orthant. Consequently, vector optimization problems (VOPs) have garnered significant attention in recent years.
Over the past two decades, descent methods have received increasing attention within the multiobjective optimization community, primarily due to the seminal work on the steepest descent method proposed by Fliege & Svaiter [17]. Inspired by Fliege and Svaiter’s contributions, researchers have extended other numerical algorithms to solve multiobjective optimization problems (MOPs) [see, e.g., 16, 40, 38, 18, 5, 36, 37, 43]. To the best of our knowledge, the study of descent methods for unconstrained vector optimization problems can be traced back to the work of Graña Drummond & Svaiter [14], who extended the steepest descent method for MOPs (SDMO) proposed by Fliege & Svaiter [17] to VOPs. In this context, the direction-finding subproblem at $x^{k}$ is formulated as follows:
@@
In recent years, complexity analysis of descent methods for MOPs has been extensively studied. Fliege et al. [19] established the convergence rates of SDMO under different convexity assumptions. Tanabe et al. [44] developed convergence results for the multiobjective proximal gradient method. Additionally, Lapucci [31] studied the complexity of a wide class of multiobjective descent methods with nonconvex assumption. However, Chen et al. [8] noted that both theoretical and empirical results indicate that existing multiobjective first-order methods exhibit slow convergence due to objective imbalances. To address this challenge, Chen et al. [7] proposed a Barzilai-Borwein descent method for MOPs (BBDMO) that dynamically tunes gradient magnitudes using Barzilai-Borwein’s rule [3] in direction-finding subproblem. From a theoretical perspective, an improved linear convergence rate is confirmed by Chen et al. [8], demonstrating that Barzilai-Borwein descent methods can effectively mitigate objective imbalances. Despite the extensive study of the complexity analysis of descent methods for MOPs, corresponding results for VOPs have received little attention. As described by Chen et al. [6], the linear convergence rates of first-order descent methods for VOPs are influenced by $C$, the base of $K^{*}$ in direction-finding subproblem. This naturally leads to the following question:
(Q1) How to select an appropriate base to accelerate convergence ?
Similarly, for the subproblem of SDMO:
@@
where $\Delta_{m}:=\{\lambda\succeq 0:\sum_{i=1}^{m}\lambda_{i}=1\}$ is a base of +$\mathbb{R}^{m}_{+}$, a natural question arises:
(Q2) Is $\Delta_{m}$ a good choice of base for +$\mathbb{R}^{m}_{+}$ in SDMO? If not, how to choose a better one?
@@
The majorization-minimization principle is a versatile tool for designing novel algorithms, which has been successful employed in nonconvex optimization [30], constrained optimization [29], and incremental optimization [34, 27]. The core idea behind the majorization-minimization method is to iteratively minimize a sequence of easily solvable upper bounds of the objective function. In this paper, we extend majorization-minimization method to vector optimization, addressing the aforementioned questions. Our work aims to provide a unified framework and convergence analysis of descent methods for VOPs from a majorization-minimization perspective. The primary contributions of this paper are summarized as follows:
@@
(i) Before presenting the majorization-minimization method and its convergence analysis, we first define the concepts of strong convexity and smoothness for a vector-valued function with respect to a partial order. Leveraging these properties, we extend the notion of the condition number to VOPs, which plays a pivotal role in establishing the linear convergence of first-order methods for VOPs.
@@
(ii) We devise a generic majorization-minimization descent method for VOPs and develop its convergence analysis. By selecting different surrogate functions, the general method can be reduced to several existing descent methods. We emphasize that the gap between the surrogate and objective functions significantly affects the performance of descent methods, which plays a central role in majorization-minimization optimization. Specifically, the steepest descent method for VOPs exhibits slow convergence due to the large gap between the surrogate and objective functions. To address this issue, we develop an improved descent method with a tighter surrogate function, resulting in improved linear convergence, and the rate of convergence is determined by the condition number. Interestingly, we show that selecting a tighter surrogate function is equivalent to using an appropriate base in the direction-find subproblem (see Remark 4.4). This provides a positive answer to (Q1). Especially, theoretical results suggest a tighter surrogate function by using Barzilai-Borwein method.
@@
(iii) We propose a Barzilai-Borwein descent method for VOPs (BBDVO) with polyhedral cones. By reformulating the subproblem, we observe that BBDMO is essentially the SDMO with an appropriately chosen base in the direction-finding subproblem (see Remark 6.3). This observation provides a positive answer to (Q2) and offers new insights into BBDMO from a majorization-minimization perspective. Notably, a VOP with a polyhedral cone can be transformed into an MOP using a transform matrix, which is often used to define a specific polyhedral cone. We demonstrate that the performance of BBDVO is insensitive to the choice of the transform matrix, whereas the steepest descent method is highly sensitive to it. Numerical results also confirm that BBDVO is an efficient method for solving VOPs with polyhedral cones.
(iv) The fast convergence of Newton method for MOPs [16] comes at a computationally expensive per-step cost. By selecting an appropriate base and reformulating the subproblem in the Newton method, we propose an efficient second-order method for MOPs that coincides with the preconditioned Barzilai-Borwein method for MOPs (PBBMO) [9]. This formulation offers novel insights into PBBMO and highlights the indispensability of the Barzilai-Borwein rule in the quasi-Newton method from a majorization-minimization perspective.
@@
The remainder of this paper is organized as follows. Section 2 introduces the necessary notations and definitions for later use. In Section 3, we present a generic majorization-minimization descent method for VOPs and analyze its convergence rates under various convexity assumptions. Sections 4 and 5 explore the connections between different descent methods from the perspective of majorization-minimization. Section 6 proposes a Barzilai-Borwein descent method specifically designed for VOPs with a polyhedral cone. In Section 7, we study second-order methods from a majorization-minimization perspective. Section 8 provides numerical results to demonstrate the efficiency of the BBDVO. Finally, conclusions are presented at the end of the paper.
@@ 2 Preliminaries
[…]
@@ 2.1 Vector optimization
[25] A vector $x^{*}\in\mathbb{R}^{n}$ is called efficient solution to (VOP) if there exists no $x\in\mathbb{R}^{n}$ such that $F(x)\preceq_{K}F(x^{\ast})$ and $F(x)\neq F(x^{\ast})$.
@@
[25] A vector $x^{*}\in\mathbb{R}^{n}$ is called weakly efficient solution to (VOP) if there exists no $x\in\mathbb{R}^{n}$ such that $F(x)\prec_{K}F(x^{\ast})$.
@@
[25] The objective function $F(\cdot)$ is called $K$-convex if
$F(\lambda x+(1-\lambda)y)\preceq_{K}\lambda F(x)+(1-\lambda)F(y),\ \forall x,y\in\mathbb{R}^{n},\ \lambda\in[0,1].$
@@
$JF(x)(y-x)\preceq_{K}F(y)-F(x),\ \forall x,y\in\mathbb{R}^{n}.$
@@ 2.2 Strong convexity and smoothness
[25] Assume that the objective function $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $x^{*}\in\mathbb{R}^{n}$ is a $K$-stationary point of (VOP) if and only if $x^{*}$ is an efficient solution of (VOP).
@@
Comparing with the smoothness and strong convexity in [6, Definitions 7 and 8] with Euclidean distance, i.e., $\omega(\cdot)=\frac{1}{2}\left\lVert{\cdot}\right\rVert$, Definitions 2.6 and 2.7 are tighter and do not depend on the reference vector $e$.
Next, we characterize the properties of the difference of two functions.
@@
    if $G(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$ and $F(\cdot)$ is $K$-smooth with $\bm{\ell}\in{\rm int}(K)$, where $\bm{\ell}\preceq\bm{\mu}$, then $H(\cdot)$ is strongly $K$-convex with $\bm{\mu}-\bm{\ell}$;
@@
Without loss of generality, we assume that
$\frac{\bm{\ell}_{1}}{\bm{\mu}_{1}}\geq\frac{\bm{\ell}_{2}}{\bm{\mu}_{2}}\geq\cdots\geq\frac{\bm{\ell}_{m}}{\bm{\mu}_{m}}.$ (2)
Then, we infer that
$\max\limits_{\lambda\in\Delta_{m}}\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}\geq\frac{\bm{\ell}_{1}}{\bm{\mu}_{1}}=\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}.$
In the following, we need to prove
$\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}\geq\max\limits_{\lambda\in\Delta_{m}}\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}.$
It is sufficient to prove that
$\frac{\bm{\ell}_{1}}{\bm{\mu}_{1}}\geq\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}$
holds for all $\lambda\in\Delta_{m}.$ Let us first state the fact that for any $a_{1}/b_{1}\geq a_{2}/b_{2}>0,c,d>0$, we have
$\frac{a_{1}}{b_{1}}\geq\frac{ca_{1}+da_{2}}{cb_{1}+db_{2}}\geq\frac{a_{2}}{b_{2}}.$ (3)
For any $\lambda\in\Delta_{m}$, we denote $\mathcal{I}(\lambda):=\{i:\lambda_{i}>0\}$ with element $i_{1}<i_{2}<\cdots<i_{|\mathcal{I}(\lambda)|}$. Substituting $a_{1}/b_{1}={\bm{\ell}_{i_{1}}}/{\bm{\mu}_{i_{1}}}$, $a_{2}/b_{2}={\bm{\ell}_{i_{2}}}/{\bm{\mu}_{i_{2}}}$, $c=\lambda_{i_{1}}$ and $d=\lambda_{i_{2}}$ into (3) yields
$\frac{\bm{\ell}_{i_{1}}}{\bm{\mu}_{i_{1}}}\geq\frac{\lambda_{i_{1}}\bm{\ell}_{i_{1}}+\lambda_{i_{2}}\bm{\ell}_{i_{2}}}{\lambda_{i_{1}}\bm{\mu}_{i_{1}}+\lambda_{i_{2}}\bm{\mu}_{i_{2}}}\geq\frac{\bm{\ell}_{i_{2}}}{\bm{\mu}_{i_{2}}}.$
For $2\leq j\leq|\mathcal{I}(\lambda)|$, by substituting $a_{1}/b_{1}={\sum_{s\in[j-1]}\lambda_{i_{s}}\bm{\ell}_{i_{s}}}/{\sum_{s\in[j-1]}\lambda_{i_{s}}\bm{\mu}_{i_{s}}}$, $a_{2}/b_{2}={\bm{\ell}_{i_{j}}}/{\bm{\mu}_{i_{j}}}$, $c=1$ and $d=\lambda_{i_{j}}$ into (3) recursively, we have
$\frac{\bm{\ell}_{i_{1}}}{\bm{\mu}_{i_{1}}}\geq\cdots\geq\frac{\sum_{s\in[j-1]}\lambda_{i_{s}}\bm{\ell}_{i_{s}}}{\sum_{s\in[j-1]}\lambda_{i_{s}}\bm{\mu}_{i_{s}}}\geq\cdots\geq\frac{\sum_{s\in[|\mathcal{I}(\lambda)|]}\lambda_{i_{s}}\bm{\ell}_{i_{s}}}{\sum_{s\in[|\mathcal{I}(\lambda)|]}\lambda_{i_{s}}\bm{\mu}_{i_{s}}}.$
This, together with the definition of $\mathcal{I}(\lambda)$ and the relation (2), implies
$\frac{\bm{\ell}_{1}}{\bm{\mu}_{1}}\geq\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}.$
Since $\lambda$ is chosen arbitrary, the proof is complete. ∎
@@
If $K=\mathbb{R}^{m}_{+}$, then $\kappa_{F,\preceq_{K}}=\max_{\lambda\in\Delta_{m}}{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}/{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}$. By Proposition 2.1, it follows that $\kappa_{F,\preceq_{K}}=\max_{i\in[m]}{\bm{\ell}_{i}}/{\bm{\mu}_{i}}$. The condition number reduces to that of MOPs [8].
@@
Assume that $F(\cdot)$ is strongly $K_{1}$-convex with $\bm{\mu}_{1}\in{\rm int}(K_{1})$ and $L_{1}$-smooth with $\bm{\ell}_{1}\in{\rm int}(K_{1})$. Then, for any pointed closed cone $K_{1}$ satisfied $K_{1}\subset K_{2}$, we have $F(\cdot)$ is strongly $K_{2}$-convex with $\bm{\mu}_{2}\in{\rm int}(K_{2})$ and $K_{2}$-smooth with $\bm{\ell}_{2}\in{\rm int}(K_{2})$. Futhermore, $\bm{\ell}_{2}\preceq_{K_{2}}\bm{\ell}_{1}$, $\bm{\mu}_{1}\preceq_{K_{2}}\bm{\mu}_{2}$, and $\kappa_{F,\preceq_{K_{2}}}\leq\kappa_{F,\preceq_{K_{1}}}$.
@@
Proposition 2.2 demonstrates that for a given vector-valued function, enlarging the order cone simplifies the corresponding vector optimization problem.
3 Majorization-minimization optimization for VOPs

### v3

# First-order Methods for Unconstrained Vector Optimization Problems: A Unified Majorization-Minimization Perspective

@@ Abstract
In this paper, we develop a unified majorization-minimization scheme and convergence analysis with first-order surrogate functions for unconstrained vector optimization problems (VOPs). By selecting different surrogate functions, the unified method can be reduced to various existing first-order methods. The unified convergence analysis reveals that the slow convergence of the steepest descent method is primarily attributed to the significant gap between the surrogate and objective functions. Consequently, narrowing this surrogate gap can enhance the performance of first-order methods for VOPs. To strike a better trade-off in terms of surrogate gap and per-iteration cost, we reformulate the direction-finding subproblem and elucidate that selecting a tighter surrogate function is equivalent to using an appropriate base of the dual cone in the direction-finding subproblem. Building on this insight, we employ the Barzilai-Borwein method to narrow the surrogate gap and propose a Barzilai-Borwein descent method for VOPs (BBDVO) with polyhedral cones. By reformulating the corresponding subproblem, we provide a novel perspective on the Barzilai-Borwein descent method, bridging the gap between this method and the steepest descent method. Finally, several numerical experiments are presented to validate the efficiency of the BBDVO.
@@ 1 Introduction
This paper focuses on the following unconstrained vector optimization problem:
@@
Let $K^{*}=\{c^{*}\in\mathbb{R}^{m}:\langle c^{*},y\rangle\geq 0,\forall y\in K\}$ be the positive polar cone of $K$, and $C$ be a compact base of $K^{*}$, namely, $C$ is a convex set such that $0\notin\text{cl}C$ and $\mathrm{cone}(C)=K^{*}$. In vector optimization, it is often impossible to improve all objectives simultaneously with respect to the partial order. Therefore, the concept of optimality is defined as efficiency [25], meaning that there is no better solution for an efficient solution. Specifically, the problem (VOP) corresponds to a multiobjective optimization problem when $K=\mathbb{R}^{m}_{+}$, where $\mathbb{R}^{m}_{+}$ denotes the non-negative orthant of $\mathbb{R}^{m}$. Various applications of multiobjective optimization problems (MOPs) can be found in engineering [35], economics [46, 18], management science [13], environmental analysis [32], machine learning [42, 48], etc. Although many real-world problems reformulated as vector-valued problems adhere to the partial order induced by $\mathbb{R}^{m}_{+}$, some applications, such as portfolio selection in securities markets [1, 2], require partial orders induced by closed convex cones other than the non-negative orthant. Consequently, vector optimization problems (VOPs) have garnered significant attention in recent years.
Over the past two decades, descent methods have received increasing attention within the multiobjective optimization community, primarily due to the seminal work on the steepest descent method proposed by Fliege and Svaiter [15]. Inspired by Fliege and Svaiter’s contributions, researchers have extended other numerical algorithms to solve multiobjective optimization problems (MOPs) [see, e.g., 14, 40, 38, 17, 5, 36, 37, 44]. To the best of our knowledge, the study of descent methods for unconstrained vector optimization problems can be traced back to the work of Graña Drummond and Svaiter [22], who extended the steepest descent method for MOPs (SDMO) proposed by Fliege and Svaiter [15] to VOPs. In this context, the direction-finding subproblem at $x^{k}$ is formulated as follows:
@@
In recent years, complexity analysis of descent methods for MOPs has been extensively studied. Fliege et al. [16] and Zeng et al. [49] established the convergence rates of SDMO under different convexity assumptions. Tanabe et al. [45] developed convergence results for the multiobjective proximal gradient method. Additionally, Lapucci [31] studied the complexity of a wide class of multiobjective descent methods with nonconvex assumption. However, Chen et al. [8] noted that both theoretical and empirical results indicate that existing multiobjective first-order methods exhibit slow convergence due to objective imbalances. To address this challenge, Chen et al. [7] proposed a Barzilai-Borwein descent method for MOPs (BBDMO) that dynamically tunes gradient magnitudes using Barzilai-Borwein’s rule [3] in direction-finding subproblem. From a theoretical perspective, an improved linear convergence rate is confirmed by Chen et al. [8], demonstrating that Barzilai-Borwein descent methods can effectively mitigate objective imbalances.
Despite the extensive study of complexity analysis for MOPs, corresponding results have received little attention in VOPs. As described by Chen et al. [6], the linear convergence rates of first-order descent methods for VOPs are essentially affected by $C$, which represents the base of $K^{*}$ in direction-finding subproblem. In general, $K^{*}$ admits infinitely many possible bases, and the choice of $C$ is therefore not unique. Although this choice critically influences the search direction and ultimately the convergence behavior of the algorithm, the existing literature offers no comprehensive theoretical analysis or guiding principle for selecting such a base. In the classical setting of multiobjective optimization, the subproblem of SDMO can be reformulated as
@@
where $\Delta_{m}:=\{\lambda\succeq 0:\sum_{i=1}^{m}\lambda_{i}=1\}$ is a base of $\mathbb{R}^{m}_{+}$. While this choice appears natural and convenient from a computational perspective, its theoretical justification and potential advantages over other alternatives remain largely unexplored. Hence, identifying an appropriate and theoretically sound base for $K^{*}$ constitutes a fundamental yet open problem, which is of both theoretical significance and practical relevance in the design of efficient first-order methods for VOPs. This naturally leads to the following question:
Can we provide a theoretical guidance for the choice of the base? (Q)
@@
To answer the proposed questions, we develop a unified framework and convergence analysis of first-order methods for VOPs from a majorization-minimization perspective. The majorization-minimization principle is a versatile tool for designing novel algorithms, which has been successful employed in nonconvex optimization [30], constrained optimization [29], and incremental optimization [34, 27]. The core idea behind the majorization-minimization method is to minimize a difficult optimization problem by iteratively minimizing a simpler surrogate function that majorizes (upper-bounds) the original objective. In this paper, we extend majorization-minimization method to vector optimization, addressing the aforementioned questions. Our work aims to provide a unified framework and convergence analysis of first-order methods for VOPs from a majorization-minimization perspective. The primary contributions of this paper are summarized as follows:
@@
(i) Before presenting the majorization-minimization method and its convergence analysis, we first define the concepts of strong convexity and smoothness for a vector-valued function with respect to a partial order. Leveraging these properties, we extend the notion of the condition number to VOPs, which plays a pivotal role in establishing the linear convergence of first-order methods for VOPs. To the best of our knowledge, it is the first definition of condition number to VOPs.
@@
(ii) We devise a unified majorization-minimization descent method for VOPs and develop its convergence analysis. By selecting different surrogate functions, the unified method can be reduced to several existing first-order methods. It is worth noting that the gap between the surrogate and objective functions significantly affects the performance of descent methods, which plays a central role in majorization-minimization optimization. Specifically, the steepest descent method for VOPs exhibits slow convergence due to the large gap between the surrogate and objective functions. To address this issue, we develop an improved descent method with a tighter surrogate function, resulting in improved linear convergence, and the rate of convergence is determined by the condition number. Interestingly, we show that selecting a tighter surrogate function is equivalent to using an appropriate base in the direction-find subproblem (see Remarks 4.4 and 5.3). This provides a positive answer to the proposed question.
@@
(iii) Theoretical results suggest a tighter surrogate function by using Barzilai-Borwein method, which motivates us to devise a Barzilai-Borwein descent method for VOPs (BBDVO) with polyhedral cones. By reformulating the subproblem, we observe that BBDVO is essentially the steepest descent method with an appropriately chosen base in the direction-finding subproblem (see Remark 6.3). Furthermore, a VOP with a polyhedral cone can be transformed into an MOP using a transform matrix, which is often used to define a specific polyhedral cone. We demonstrate that the performance of BBDVO is insensitive to the choice of the transform matrix, whereas the steepest descent method is highly sensitive to it. From a majorization–minimization perspective, we further elucidate why the line search procedure in VOPs leads to a slower linear convergence rate compared with its counterpart without line search. In contrast, the backtracking strategy preserves the same linear convergence rate as first-order methods for VOPs without line search.
@@
The remainder of this paper is organized as follows. Section 2 introduces the necessary notations and definitions for later use. In Section 3, we present a generic majorization-minimization descent method for VOPs and analyze its convergence rates under various convexity assumptions. Sections 4 and 5 explore the connections between different descent methods from the perspective of majorization-minimization. Section 6 proposes a Barzilai-Borwein descent method for VOPs with a polyhedral cone. Section 7 provides numerical results to demonstrate the efficiency of the BBDVO. Finally, conclusions are presented at the end of the paper.
@@ 2 Preliminaries
Throughout this paper, $\mathbb{R}^{n}$ and $\mathbb{R}^{m\times n}$ denote the set of $n$-dimensional real column vectors and the set of $m\times n$ real matrices, respectively. The space $\mathbb{R}^{n}$ is equipped with the inner product $\langle\cdot,\cdot\rangle$ and the induced norm $\|\cdot\|$. The interior, boundary and the closure of a set are denoted by $\mathrm{int}(\cdot)$, $\mathrm{bd}(\cdot)$ and $\mathrm{cl}(\cdot)$, respectively. The cone generated by a set is denoted by $\textrm{cone}(\cdot)$. For simplicity, we denote $m]:=\{1,2,...,m\}$, $\mathbf{1}_{m}$ and $I_{m}$ the all-ones vector in $\mathbb{R}^{m}$ and identity matrix in $\mathbb{R}^{m\times m}$, respectively.
Since $K$ is a closed convex cone, it follows that $K=K^{**}$ (see [41, Theorem 14.1]),
$K=\{y\in\mathbb{R}^{m}:\langle y,c^{*}\rangle\geq 0,\forall c^{*}\in K^{*}\},$
and
$\textrm{int}(K)=\{y\in\mathbb{R}^{m}:\langle y,c^{*}\rangle>0,\forall c^{*}\in K^{*}\setminus\{0\}\}.$
Since $\textrm{int}(K)\neq\emptyset$, we assume that there exists a compact and convex set $C$ such that
$0$ $\notin C,$ (1)
$\operatorname{cone}(C)$ $=K^{*}.$ (2)
Therefore
$K=\{y\in\mathbb{R}^{m}:\langle y,c^{*}\rangle\geq 0,\forall c^{*}\in C\},$ (3)
$\textrm{int}(K)=\{y\in\mathbb{R}^{m}:\langle y,c^{*}\rangle>0,\forall c^{*}\in C\}.$ (4)
The latter equality, together with the compactness of $C$, implies that
$\min_{c^{*}\in C}\{\left\langle{c^{*},y}\right\rangle\}>0,~~~\forall y\in\mathrm{int}(K).$ (5)
For more details on $C$, we refer the readers to [22, pp. 400].
@@ 2.1 Vector optimization
[25, Definition 11.3] A vector $x^{*}\in\mathbb{R}^{n}$ is called efficient solution to (VOP) if there exists no $x\in\mathbb{R}^{n}$ such that $F(x)\preceq_{K}F(x^{\ast})$ and $F(x)\neq F(x^{\ast})$.
@@
[25, Definition 11.5] A vector $x^{*}\in\mathbb{R}^{n}$ is called weakly efficient solution to (VOP) if there exists no $x\in\mathbb{R}^{n}$ such that $F(x)\prec_{K}F(x^{\ast})$.
@@
[25, Definition 2.4] The objective function $F(\cdot)$ is called $K$-convex if
$F(\lambda x+(1-\lambda)y)\preceq_{K}\lambda F(x)+(1-\lambda)F(y)$
holds for all $x,y\in\mathbb{R}^{n},\ \lambda\in[0,1]$.
@@
$JF(x)(y-x)\preceq_{K}F(y)-F(x)$
holds for all $x,y\in\mathbb{R}^{n}$. (see [25, Theorem 2.20]).
@@ 2.2 Strong convexity and smoothness
[14, Theorem 3.1] Assume that the objective function $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $x^{*}\in\mathbb{R}^{n}$ is a $K$-stationary point of (VOP) if and only if $x^{*}$ is an efficient solution of (VOP).
@@
Comparing with the smoothness and strong convexity in [6, Definitions 7 and 8] with Euclidean distance, i.e., $\omega(\cdot)=\frac{1}{2}\left\lVert{\cdot}\right\rVert^{2}$, Definitions 2.6 and 2.7 are tighter and do not depend on the reference vector $e$.
Next, we characterize the properties of the difference of two vector-valued functions.
@@
[…]
@@
Since $\bm{\ell},\bm{\mu}\in\mathbb{R}^{m}_{++}$, for any $i\in[m]$, we have
$\bm{\ell}_{i}\leq\bm{\mu}_{i}\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}.$
Multiply by $\lambda_{i}\geq 0$ and sum over $i\in[m]$:
$\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}\leq\left(\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}\right)\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}.$
Dividing by the positive $\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}$ yields
$\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}\leq\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}.$
Therefore, the relation
$\max\limits_{\lambda\in\Delta_{m}}\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}\leq\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}$
holds due to the arbitrary of $\lambda$. Let $s$ be an index where the maximum ratio is attained, i.e.
$\frac{\bm{\ell}_{s}}{\bm{\mu}_{s}}=\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}.$
Take $\lambda_{s}=1$ and $\lambda_{i}=0$ for $i\neq s$, we have
$\max\limits_{\lambda\in\Delta_{m}}\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}\geq\frac{\bm{\ell}_{s}}{\bm{\mu}_{s}}=\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}.$
This completes the proof. ∎
@@
If $K=\mathbb{R}^{m}_{+}$, then $\kappa_{F,\preceq_{K}}=\max_{\lambda\in\Delta_{m}}{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}/{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}$. By Proposition 2.1, it follows that $\kappa_{F,\preceq_{K}}=\max_{i\in[m]}{\bm{\ell}_{i}}/{\bm{\mu}_{i}}$. In other words, the condition number for multiobjective optimization is the largest condition number among all objective functions.
@@
Assume that $F(\cdot)$ is strongly $K_{1}$-convex with $\bm{\mu}_{1}\in{\rm int}(K_{1})$ and $K_{1}$-smooth with $\bm{\ell}_{1}\in{\rm int}(K_{1})$. Then, for any order cone $K_{2}$ satisfied $K_{1}\subset K_{2}$, we have $F(\cdot)$ is strongly $K_{2}$-convex with $\bm{\mu}_{2}\in{\rm int}(K_{2})$ and $K_{2}$-smooth with $\bm{\ell}_{2}\in{\rm int}(K_{2})$. Futhermore, $\bm{\ell}_{2}\preceq_{K_{2}}\bm{\ell}_{1}$, $\bm{\mu}_{1}\preceq_{K_{2}}\bm{\mu}_{2}$, and $\kappa_{F,\preceq_{K_{2}}}\leq\kappa_{F,\preceq_{K_{1}}}$.
@@
Proposition 2.2 shows that, for a fixed vector-valued function, enlarging the underlying order cone effectively simplifies the associated vector optimization problem; see Lemma 4.2 for a formal statement. This observation motivates the use of a larger order cone to accelerate first-order methods for VOPs.
3 Majorization-minimization with first-order surrogate functions for VOPs